OpenStudy (anonymous):

Let f(x) = x2 - 3x - 7. Find f(-3). can someone walk me through this?

OpenStudy (solomonzelman):

plug in -3 for x

OpenStudy (calculusxy):

\[f(-3)=-3^2 - 7\] Solve the equation

OpenStudy (solomonzelman):

you are forgetting something, also don't do someone else's work.

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[f(x) = x^2 - 3x - 7\] \[f(\spadesuit )=\spadesuit ^2-3\times \spadesuit -7\]

OpenStudy (anonymous):

\[f(\heartsuit )=\heartsuit ^2-3\times \heartsuit -7\]

OpenStudy (anonymous):

\[f(-3)=(-3)^2-3\times (-3)-7\]